OpenStudy (anonymous):

What is the solution to the system of inequalities: x – 2y > –16 3x + 4y < 20

OpenStudy (anonymous):

Sketch the graph of each line in the plane; x-2y=-16 has intercepts (0,8) and (-16, 0) and you would shade below the line 3x+4y=20 has intercepts (0,5) and (20/3, 0) and you would shade below the line I made an error with the sketch; gimme a minute to do it again)

OpenStudy (anonymous):

here is the sketch

OpenStudy (anonymous):

x-2y>-16 so x=2y-16 plug this x in for second equation 3(2y-16)+4y<20 now solve this for y 10y-48<20 so y is 68/10=34/5 now solve for x using x=2y-16 x=2(34/5)-16=(66/5)-16=-14/5 so the solution to these is the point (-14/5,34/5)

OpenStudy (anonymous):

thts not one of the choices.. the choices are (8, 3) (11, –1) (–4, –6) (–2, 7)

OpenStudy (anonymous):

I really need help..thank you soo much you guyzz

OpenStudy (anonymous):

hey katrina

OpenStudy (katrinakaif):

Hello :)

OpenStudy (anonymous):

MeEngr, thnks for the help...but are you sure thats the right answer??

OpenStudy (anonymous):

u know...you can work backwards and check if any of the choices make the question true...so for example...for x=8 y=3 plug in the system and check if they are true so 2>-16 and 36<20 no thats not true so choice number one is not right...move on to next one x=11 y=-1 so 13>-16 and 29<20 so thats not right..next one x=-4 y=-6 so 8>-16 and -36<20 so thats true..last one x=-1 y=7 so 13>-15 and 25<20 not true so only solution true is (-4,-6) because it satisfies the inequalities now